# Planning and Rendering: Towards End-to-End Product Poster Generation

### v1

@@ Abstract
End-to-end product poster generation significantly optimizes design efficiency and reduces production costs. Prevailing methods predominantly rely on image-inpainting methods to generate clean background images for given products. Subsequently, poster layout generation methods are employed to produce corresponding layout results. However, the background images may not be suitable for accommodating textual content due to their complexity, and the fixed location of products limits the diversity of layout results. To alleviate these issues, we propose a novel product poster generation framework named P&R. The P&R draws inspiration from the workflow of designers in creating posters, which consists of two stages: Planning and Rendering. At the planning stage, we propose a PlanNet to generate the layout of the product and other visual components considering both the appearance features of the product and semantic features of the text, which improves the diversity and rationality of the layouts. At the rendering stage, we propose a RenderNet to generate the background for the product while considering the generated layout, where a spatial fusion module is introduced to fuse the layout of different visual components. To foster the advancement of this field, we propose the first end-to-end product poster generation dataset PPG30k, comprising 30k exquisite product poster images along with comprehensive image and text annotations. Our method outperforms the state-of-the-art product poster generation methods on PPG30k. The PPG30k will be released soon.
[…]
@@ 1 Introduction
Product posters are important for product promotion. An exquisite poster should not only have a rational layout of visual components, such as underlays, texts, and the product but also have a pleasant background. This challenging task is typically undertaken by professional designers, which leads to low efficiency and high cost in poster production. To generate high-quality posters with low cost, we propose a new task, End-to-end product Poster Generation (EPG), which aims to generate an image that transmits the product messages to social groups given the product image and texts.
Figure 1: Two fields related to EPG: (a) Image-inpainting and (b) Poster layout generation. (c) The combination of (a) and (b) can be regarded as a naive solution to EPG. (d) The workflow of professional designers. “score” refers to the confidence of layout prediction, and the gray blocks are translations of texts. Green and red boxes represent texts and underlays, respectively.
The most related fields to EPG are image-inpainting [20, 30, 40, 26, 1, 41, 31] and poster layout generation [44, 36, 13, 21]. As Figure 1 (a) and (b) show, the image-inpainting model completes the background of the poster based on the product image, and the poster layout generator plans the layout of visual components given a manually designed image. As Figure 1 (c) shows, simply combining the two methods can be regarded as a naive solution to EPG, where the result of the image-inpainting model can be put into the layout generator to arrange the visual components. However, this solution still has the following issues: 1) The background generated from the image-inpainting model may be too complex to display any text, which leads to a higher failure rate of layout prediction. 2) The image-inpainting methods pre-fix the location of the product, so that the layout generator can only arrange the locations of texts and underlays. This limits the diversity of poster design.
@@
Referring to the designing process of professional designers, we put forward the first product poster generation framework P&R shown as Figure 2. At the planning stage, we propose the PlanNet to generate layouts of products and other visual components. PlanNet firstly encodes the product image and the texts, and then fuses them with a layout decoder to predict more reasonable and flexible layouts. At the rendering stage, we propose the RenderNet to integrate layouts and the product image into control conditions, where a spatial fusion module is used to explore the spatial relationships among layouts generated by PlanNet, and the appearance of the product is also encoded to ensure harmony between the generated background and the product. Finally, these two control conditions are input to ControlNet [41], which can guide Stable Diffusion [31] to generate product posters with pleasant backgrounds and easy to convey information.
To train and evaluate our method, we construct the first product poster generation dataset PPG30k including appealing product posters, layouts of visual components, product masks, and text contents. We conduct comprehensive experiments on PPG30k, and the results indicate that our method performs better in terms of the rationality of the layout and the aesthetics of the image. The contributions of our paper are four-fold:
  • We propose a new task EPG and the first product poster generation framework P&R, which can generate appealing posters to transmit product messages effectively.
@@
  • We build the first large-scale dataset PPG30k for training and evaluating the models of EPG.
@@ 2.2 Conditional Image Generation
The conditional generation models aim to generate images according to the given conditions. At the rendering stage of our poster generating system, the generator should consider the product image and the layout of visual components, which is related to the research of image-inpainting [20, 30, 40, 26, 1, 41, 31], and layout-to-image [11, 5, 6, 42, 38, 37, 41].
@@ 3 Method
The framework of P&R is shown as Figure 2, which consists of a PlanNet and a RenderNet. The PlanNet predicts the overall layout of the poster based on the given product image and texts, and then the RenderNet renders the predicted layout as a poster image. For the PlanNet, it first encodes the product image and texts with the image and text encoders. Then we input the encoded features to a layout decoder to learn the relationship among layout elements, image features, and text features. After a $T_{P}$-step denoising process, the initialized layout is restored to the overall layout of the poster. For the RenderNet, it utilizes the layout results generated by PlanNet with two branches. The first layout branch fuses the layouts of visual components into a unified layout embedding with a spatial fusion module, which can learn the spatial relationship among the different visual components. The second vision branch uses a series of convolution layers to extract the visual embeddings of the repositioned product, which contributes to generating a harmonious background with the product. Then extracted layout and visual embeddings are put into the ControlNet to guide the generation process of stable diffusion. After generating an image prepared for rendering texts, we use a text-rendering module to obtain the final product poster.
@@ 3.1 PlanNet
A priori overall planning layout is beneficial for the downstream generation tasks. Although content-agnostic methods can predict the overall layout, they neglect the appearance of the product and the text content, making it difficult to achieve a good visual balance. To address this issue, we propose the PlanNet, which considers the product and text information simultaneously while predicting the overall layout. The layout consists of a set of elements, which are represented by five attributes $c,x,y,w,h$, where $c$ is the category, $(x,y)$ is the center point, and $w$, $h$ are the width and height. Each attribute has $K$ different values. To handle discrete variables, the PlanNet is based on discrete diffusion models including diffusion and denoising processes.
@@ 3.2 RenderNet
At the inference stage, the well-trained RenderNet predicts the noise added to an image for $T_{R}$ steps, given a random noise $Z_{T_{R}}$, text embedding $\tau$, the control condition $\epsilon^{c}_{\theta}(Z^{\prime})$. After removing the noise, we obtain the image prepared for placing texts.
As it is insufficient to train a visual text generation model solely on the poster image dataset, we introduce an extra text-rendering module to render texts on the generated image heuristically. We choose the appropriate text color based on the main color of the text area, and then randomly select fonts from the font library. The code will be released and we will focus on integrating a visual text generation model into the RenderNet in the future.
@@ 4 Dataset
Figure 5: Some visual examples in PPG30k.
Figure 6: (a) Poster layout annotations. Product, texts and underlay are annotated. The texts in gray boxes are the translations of the texts on the posters (b) The clean image. The texts on the poster are erased while underlays are retained.
To promote research in this field, we construct the first large-scale end-to-end product poster generation dataset PPG30k. The images are collected from CGL-Dataset [44], where we remove posters with portraits to avoid portrait rights disputes, and posters with unsightly backgrounds to improve generation quality. The selected images are with pleasing backgrounds and can effectively transmit messages, as shown in Figure 5. Eventually, there are 32724 training images and 1426 testing images in PPG30k, and the image size is $513\times 750$.
@@ 5.1 Implementation Details
For the PlanNet, there are 4 transformer blocks with 8 attention heads, 512 embedding dimensions, and 2048 hidden dimensions, and we set the dropout rate 0.1. The diffusion timesteps $T_{P}$ is 100. We train the PlanNet for 100 epochs using AdamW optimizer [25] with a learning rate of $5.0\times 10^{-4}$. For the RenderNet, we train it for 500 epochs with the same optimizer as ControlNet [41]. The prompt is uniformly set as “a poster of an e-commerce product”. We resize the image to $256\times 384$. The hidden feature map of the spatial fusion module satisfies $C=320,W=128,H=192$. The layer number $S$ is 2, the inference step $T_{R}$ is 50, and the patch size $P$ is 8.
5.2 Evaluation Metrics
User Study. To evaluate the quality of the poster, we conduct a user study on 20 experienced advertising practitioners. In this study, we use 100 groups of images, where each group contains 4 posters generated by different methods. All these results in each group are presented side-by-side and in a random order. Practitioners are asked to rank them from the perspectives of background quality (BQ) and layout rationality (LR), where 1 is the worst and 4 is the best. When ranking the BQ, the practitioners focus on the suitability of serving as the product poster and the harmony between the background and the product. When ranking the LR, the practitioners consider the overlap of different components and the visual balance of components on the poster. To evaluate the performances of poster layout generators, we use 100 groups of layouts, where each group contains 4 randomly ordered layouts generated by different methods. Practitioners are asked to judge whether the layout result is qualified (${\rm R_{pass}}$) and select the best layout (${\rm R_{best}}$). When evaluating the two metrics, practitioners consider the overlap of different components and, rationality of visual components.
5.3 Quantitative Analysis
@@
Comparison with SOTA Poster Generators. As mentioned above, combining image-inpainting and poster layout generation methods can be seen as a naive solution to EPG. To verify the superiority of P&R, we combine three image-inpainting methods with a SOTA layout generator CGL-GAN [44] and compare their generated posters with ours.
As Table 1 shows, our method performs the best on FID-image and CLIP-score, which indicates that the style of background images generated by our RenderNet is closer to that of product posters. We also calculate the average BQ and LR. The results show that the images generated by our method are more suitable to serve as product posters, and our generated layouts achieve a better visual balance.
We take the SD-inpaint method as an example to compare our method with SOTA poster generators qualitatively. As Figure 7 shows, although the SD-inpaint method can generate realistic background images, the content of the background images is too complex to have sufficient space for visual components, which results in a high failure rate of layout generation. Our method can generate an appealing background with the superiority of transmitting product information.
Comparison with SOTA Layout Generators. To verify the advantages of our PlanNet, we compare the PlanNet with SOTA layout generators, including LayoutDM [16], FlexDM [15] and BoxNet [34]. Based on their official codes, we train these models on PPG30k. The comparison result is shown in Table 2. Because of making fuller use of text and image features, our method achieves the best on FID-layout and Max IoU metrics. Additionally, to verify the impact of layout on background generation, we use the RenderNet to generate background based on the layout results of four methods. The images generated from the results of the PlanNet achieve the best on FID-image and CLIP-score metrics, which proves that our PlanNet is more beneficial for downstream image generation. Meanwhile, the result of the user study indicates that our PlanNet arranges various elements with better composition.
Comparison with SOTA Conditional Generators. Layout-to-image methods ignore the product content when generating posters, resulting in the generated background may not match the style of the product. Our RenderNet combines the advantages of both image-inpainting and layout-to-image methods, which can give attention to both layouts and the product. To verify this, we compare two state-of-the-art layout-to-image methods based on the ground truth layouts. Table 3 indicates that backgrounds generated by our method match the product more harmoniously.
@@ 5.4 Qualitative Analysis
Diversity. To show the diversity of P&R, we set different random seeds to generate posters. As shown in Figure 8, the leftmost column is the input product and texts. In the rest part, the first row is the different layouts and the second row is the posters generated based on the layouts above. This shows that our method can generate a variety of reasonable layouts and appealing posters.

### v2

@@ Abstract
Product poster generation significantly optimizes design efficiency and reduces production costs. Prevailing methods predominantly rely on image-inpainting methods to generate clean background images for given products. Subsequently, poster layout generation methods are employed to produce corresponding layout results. However, the background images may not be suitable for accommodating textual content due to their complexity, and the fixed location of products limits the diversity of layout results. To alleviate these issues, we propose a novel product poster generation framework based on diffusion models named P&R. The P&R draws inspiration from the workflow of designers in creating posters, which consists of two stages: Planning and Rendering. At the planning stage, we propose a PlanNet to generate the layout of the product and other visual components considering both the appearance features of the product and semantic features of the text, which improves the diversity and rationality of the layouts. At the rendering stage, we propose a RenderNet to generate the background for the product while considering the generated layout, where a spatial fusion module is introduced to fuse the layout of different visual components. To foster the advancement of this field, we propose the first product poster generation dataset PPG30k, comprising 30k exquisite product poster images along with comprehensive image and text annotations. Our method outperforms the state-of-the-art product poster generation methods on PPG30k. The PPG30k will be released soon.
keywords:
Poster Generation, Diffusion Model, Generative Multimedia, Multimodal Fusion.
@@ 1 Introduction
Product posters are important for product promotion. An exquisite poster should not only have a rational layout of visual components, such as underlays, texts, and the product but also have a pleasant background. This challenging task is typically undertaken by professional designers, which leads to low efficiency and high cost in poster production. To generate high-quality posters with low cost, we propose a new task, Product Poster Generation (PPG), which aims to generate an image that transmits the product messages to social groups given the product image and texts.
The most related fields to PPG are image-inpainting [1, 2, 3, 4, 5, 6, 7] and poster layout generation [8, 9, 10, 11]. As Figure 1 (a) and (b) show, the image-inpainting model completes the background of the poster based on the product image, and the poster layout generator plans the layout of visual components given a manually designed image. As Figure 1 (c) shows, simply combining the two methods can be regarded as a naive solution to PPG, where the result of the image-inpainting model can be put into the layout generator to arrange the visual components. However, this solution still has the following issues: 1) The background generated from the image-inpainting model may be too complex to display any text, which leads to a higher failure rate of layout prediction. 2) The image-inpainting methods pre-fix the location of the product, so that the layout generator can only arrange the locations of texts and underlays. This limits the diversity of poster design.
@@
Figure 1: Two fields related to CPG: (a) Image-inpainting and (b) Poster layout generation. (c) The combination of (a) and (b) can be regarded as a naive solution to CPG. (d) The workflow of professional designers. “score” refers to the confidence of layout prediction, and the gray blocks are translations of texts. Green and red boxes represent texts and underlays, respectively.
Referring to the designing process of professional designers, we put forward the first product poster generation framework P&R with diffusion models, shown as Figure 2. At the planning stage, we propose the PlanNet to generate layouts of products and other visual components. PlanNet firstly encodes the product image and the texts, and then fuses them with a layout decoder to predict more reasonable and flexible layouts.
At the rendering stage, we propose the RenderNet to integrate layouts and the product image into control conditions, where a spatial fusion module is used to explore the spatial relationships among layouts generated by PlanNet, and the appearance of the product is also encoded to ensure harmony between the generated background and the product.
To train and evaluate our method, we construct the first product poster generation dataset PPG30k including appealing product posters, layouts of visual components, product masks, and text contents. We conduct comprehensive experiments on PPG30k, and the results indicate that our method performs better in terms of the rationality of the layout and the aesthetics of the image.
The contributions of our paper are four-fold:
  • We propose a new task PPG and the first product poster generation framework P&R, which can generate appealing posters to transmit product messages effectively.
@@
  • We build the first large-scale dataset PPG30k for training and evaluating the models of PPG.
Our paper is organized as follows: In Section 2, we summarize the latest development in layout and conditional image generation. In Section 3, we show the details of our proposed method P&R. In Section 3.4, we introduce the construction of our dataset PPG30k. In Section 5, we compare our proposed method with SOTA methods. In the last section, we draw a conclusion and discuss the future work.
@@ 2.2 Conditional Image Generation
The conditional generation models aim to generate images and videos according to the given conditions [31, 32, 33, 34, 35]. At the rendering stage of our poster generating system, the generator should consider the product image and the layout of visual components, which is related to the research of image-inpainting [1, 2, 3, 4, 5, 6, 7, 36], and layout-to-image [37, 38, 39, 40, 41, 42, 6].
@@ 3 Method
3.1 Task Definition
Tasks Background Text Product Layout
PLG $\times$ $\times$ $\checkmark$ $\checkmark$
PG $\times$ $\checkmark$ $\checkmark$ $\times$
BG $\checkmark$ $\times$ $\checkmark$ $\times$
VTG $\checkmark$ $\checkmark$ $\times$ $\times$
PPG $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Comparison of PPG with other tasks.
Product Poster Generation (PPG) is defined as generating product poster images that contain given products and text content. Tasks similar to PPG include Poster Layout Generation (PLG) [8, 11], Poster Generation (PG) [44], Background Generation (BG) [1, 45], and Visual Text Generation (VTG) [46, 47, 48]. There are four distinct factors that make PPG more challenging compared to these tasks as outlined in Table 1: 1) Background: The background image is generated automatically rather than manually designed. 2) Text: Generated images need to include visual text with the given content. 3) Product: The identity of the product shown in posters should be exactly consistent with the given product. 4) Layout: The positions of products and texts are automatically generated rather than manually designed.
3.2 Overall Framework
The framework of P&R is shown as Figure 2, which consists of a PlanNet and a RenderNet. The PlanNet predicts the overall layout of the poster based on the given product image and texts, and then the RenderNet renders the predicted layout as a poster image. For the PlanNet, it first encodes the product image and texts with the image and text encoders. Then we input the encoded features to a layout decoder based on discrete diffusion model to learn the relationship among layout elements, image features, and text features. After a $T_{P}$-step denoising process, the initialized layout is restored to the overall layout of the poster. For the RenderNet, it utilizes the layout results generated by PlanNet with two branches. The first layout branch fuses the layouts of visual components into a unified layout embedding with a spatial fusion module, which can learn the spatial relationship among the different visual components. The second vision branch uses a series of convolution layers to extract the visual embeddings of the repositioned product, which contributes to generating a harmonious background with the product. Then extracted layout and visual embeddings are put into the ControlNet to guide the generation process of stable diffusion. After generating an image prepared for rendering texts, we use a text-rendering module to obtain the final product poster.
@@ 3.3 PlanNet
A prior overall planning layout is beneficial for the downstream generation tasks. Although content-agnostic methods can predict the overall layout, they neglect the appearance of the product and the text content, making it difficult to achieve a good visual balance. To address this issue, we propose the PlanNet, which considers the product and text information simultaneously while predicting the overall layout. The layout consists of a set of elements, which are represented by five attributes $c,x,y,w,h$, where $c$ is the category, $(x,y)$ is the center point, and $w$, $h$ are the width and height. Each attribute has $K$ different values. To handle discrete variables, the PlanNet is based on discrete diffusion models including diffusion and denoising processes.
@@ 3.4 RenderNet
At the inference stage, the well-trained RenderNet predicts the noise added to an image for $T_{R}$ steps, given a random noise $Z_{T_{R}}$, text embedding $\tau$, the control condition $\epsilon^{c}_{\theta}(Z^{\prime})$. After removing the noise, we obtain the image $Z_{0}$ prepared for placing texts.
Figure 5: Some visual examples in PPG30k.
As it is insufficient to train a visual text generation model solely on the poster image dataset, we introduce an extra text-rendering module to render texts on the generated image heuristically. Specifically, assuming that the location of the i-th text element that needs to render is ${Text}_{i}$. To determine the text color, we first filter out pixels with lower transparency within ${Text}_{i}$ of $Z_{0}$ to obtain a set of pixels that contribute more to visual perception. Then, we use the color with the highest proportion in the set as the representative color of ${Text}_{i}$. Finally, we offset the representative color in HSV space as the text color, and the offset value needs to ensure that the color tone before and after the offset is consistent, and the aesthetic matching is harmonious. To determine the text font, we collect over twenty commonly used font libraries in advance, and then randomly select one as the text font. The relevant code will be released to help reproduce the results. Although some pre-trained text generation and editing models [46, 47, 48] can also be used to render text in designated areas, we have empirically found that the generated text is prone to errors, especially for Chinese, and the heuristic rule-based text rendering module can accurately and clearly generate visual text. In the future, we will focus on improving the ability of visual text generation models and inheriting them into RenderNet.
@@ 4 Dataset
Figure 6: (a) Poster layout annotations. Product, texts and underlay are annotated. The texts in gray boxes are the translations of the texts on the posters. (b) The clean image. The texts on the poster are erased while underlays are retained.
To promote research in this field, we construct the first large-scale product poster generation dataset PPG30k. The images are collected from CGL-Dataset [8], where we remove posters with portraits to avoid portrait rights disputes, and posters with unsightly backgrounds to improve generation quality. The selected images are with pleasing backgrounds and can effectively transmit messages, as shown in Figure 5. Eventually, there are 32724 training images and 1426 testing images in PPG30k, and the image size is $513\times 750$.
@@ 5.1 Implementation Details
Since RenderNet requires precise control, and there is noise in the layouts predicted by PlanNet, which is harmful to network training. Therefore, we train PlanNet and RenderNet separately and use ground-truth layouts when training RenderNet.
@@
For the PlanNet, there are 4 transformer blocks with 8 attention heads, 512 embedding dimensions, and 2048 hidden dimensions, and we set the dropout rate 0.1. The diffusion timesteps $T_{P}$ is 100. We train the PlanNet for 100 epochs using AdamW optimizer [54] with a learning rate of $5.0\times 10^{-4}$. The training time is 12 hours on a single P40 GPU. For the RenderNet, it takes 1 day on 24 A100 GPUs to train for 500 epochs with the same optimizer as ControlNet [6]. The hidden feature map of the spatial fusion module satisfies $C=320,W=128,H=192$. The layer number $S$ is 2, the inference step $T_{R}$ is 50, and the patch size $P$ is 8.
5.2 Evaluation Metrics
@@
User Study. To evaluate the quality of the poster, we conduct a user study on 20 experienced advertising practitioners. In this study, we use 100 groups of images, where each group contains 4 posters generated by different methods. All these results in each group are presented side-by-side and in a random order. Practitioners are asked to rank them from the perspectives of background quality (BQ) and layout rationality (LR), where 1 is the worst and 4 is the best. When ranking the BQ, the practitioners focus on the suitability of serving as the product poster and the harmony between the background and the product. When ranking the LR, the practitioners consider the overlap of different components and the visual balance of components on the poster.
To avoid the influence of font on the ranking results of BQ and LR, there is no text on the poster, and all the visual components are represented by boxes. To evaluate the performances of poster layout generators, we use 100 groups of layouts, where each group contains 4 randomly ordered layouts generated by different methods. Practitioners are asked to judge whether the layout result is qualified (${\rm R_{pass}}$) and select the best layout (${\rm R_{best}}$). When evaluating the two metrics, practitioners consider the overlap of different components and, rationality of visual components.
5.3 Quantitative Analysis
Method $R_{com}\downarrow$ $R_{sub}\downarrow$ $R_{shm}$downarrow
LayoutVTN [56] 41.77 1.32 22.21
LayoutTransformer [24] 40.92 1.31 21.08
LayoutDM [29] 40.91 1.24 18.42
PlanNet 39.31 1.02 15.39
Table 4: Layout generation results on CGL-Dataset.
@@ 5.3 Quantitative Analysis
Figure 9: Density-coverage trade-off of different methods.
Comparison with SOTA Poster Generators. As mentioned above, combining image-inpainting and poster layout generation methods can be seen as a naive solution to PPG. To verify the superiority of P&R, we combine three pre-trained image-inpainting models with a SOTA layout generator CGL-GAN [8] and compare their generated posters with ours.
As Table 2 shows, our method performs the best on FID-image and CLIP-score, which indicates that the style of background images generated by our RenderNet is closer to that of product posters. We also calculate the average BQ and LR. The results show that the images generated by our method are more suitable to serve as product posters, and our generated layouts achieve a better visual balance. We take the SD-inpaint method as an example to compare our method with SOTA poster generators qualitatively. As Figure 7 shows, although the pre-trained SD-inpaint model can generate realistic background images, the content of the background images is too complex to have sufficient space for visual components, which results in a high failure rate of layout generation. Our method can generate an appealing background with the superiority of transmitting product information.
@@
Comparison with SOTA Layout Generators. To verify the advantages of our PlanNet, we compare the PlanNet with SOTA layout generators, including LayoutDM [29], FlexDM [25] and BoxNet [26]. Based on their official codes, we train these models on PPG30k. The comparison result is shown in Table 3. Because of making fuller use of text and image features, our method achieves the best on FID-layout and Max IoU metrics. Additionally, to verify the impact of layout on background generation, we use the RenderNet to generate background based on the layout results of four methods. The images generated from the results of the PlanNet achieve the best on FID-image and CLIP-score metrics, which proves that our PlanNet is more beneficial for downstream image generation. Meanwhile, the result of the user study indicates that our PlanNet arranges various elements with better composition. As shown in Table 4, we also compare the performance on public CGL-Dataset, which further proves the superiority of PlanNet.
@@
Comparison with SOTA Conditional Generators. Layout-to-image methods ignore the product content when generating posters, resulting in the generated background may not match the style of the product. Our RenderNet combines the advantages of both image-inpainting and layout-to-image methods, which can give attention to both layouts and the product. To verify this, we finetune two state-of-the-art layout-to-image methods based on the ground truth layouts. Our P&R method takes product images as additional inputs, which makes the generated background more suitable for specific products. Table 5 indicates that backgrounds generated by our method match the product more harmoniously.
@@ 5.4 Qualitative Analysis
Diversity. To show the diversity of P&R, we set different random seeds to generate posters. As shown in Figure 8, the leftmost column is the input product and texts. In the rest part, the first row is the different layouts and the second row is the posters generated based on the layouts above. This shows that our method can generate a variety of reasonable layouts and appealing posters. We also plot the density and coverage metrics [29] of PlanNet and LayoutDM in Figure 9, which proves that our method achieves a better trade-off between the diversity and fidelity.